OpenStudy (itrymath):

Choose the slope-intercept equation of the line that passes through the point shown and is perpendicular to the line shown. (4 points) y = - x y = x - 6 y = - x - y = x -

OpenStudy (itrymath):

Choose the slope-intercept equation of the line that passes through the point shown and is perpendicular to the line shown. (4 points) y = -3/2 x y = 3/2 x - 6 y = -2/3 x - y = 2/3 x - 13/3

OpenStudy (itrymath):

@Will.H

OpenStudy (itrymath):

@563blackghost

OpenStudy (will.h):

Did you found the slope of the line in the graph?

OpenStudy (itrymath):

WAIT YOU TRADE

OpenStudy (will.h):

I do

OpenStudy (itrymath):

i got 2 over 3 @Will.H

OpenStudy (itrymath):

@Will.H

OpenStudy (will.h):

Sorry for late. Let's see if the slope is right (0,3) (3,5) Slope = 2/3 Therefore the other slope that would be perpendicular to that is (2/3)x = -1 X = -1.5 They said that passes through the point... Yet u didn't type in the point

OpenStudy (will.h):

-1.5 means also -3/2

OpenStudy (will.h):

Oh so there is no point. That means they just want the slope Y = (-3/2)x Make sense?